7.) A 24.6 ml solution of unknown acid has 0.504 g/L of acid in solution. Calculate the molar mass of the solution, and use that to calculate the concentration and pH of the solution.
1
Expert's answer
2020-05-07T14:09:11-0400

Solution.

"C = \frac{m}{M \times V} = \frac{\rho}{M}"

"M = \frac{\rho}{(\frac{V}{Vm})} = 458.18 \ g/mol"

"C = 0.0011 \ M"

"pH = -lg[H^+] = 2.96"

Answer:

M = 458.18 g/mol

C = 0.0011 M

pH = 2.96